# From Landau two-fluid model to de Sitter Universe

### v9

@@ IX.8 Acoustic Planck constants
Returning to the two-fluid hydrodynamics, let us consider quasiparticles in Bose superfluid 4He. The flow of the liquid provides the acoustic gravity for ”relativistic” quasiparticles – phonons. The effective acoustic metric and thus the acoustic Planck constants, $\hbar_{\rm ac}$ and $/\!\!\!h_{\rm ac}$, can be expressed in terms of the parameters of this liquid.
@@
[…]
@@
The quantum mechanics of phonons demonstrates that it can be described by the effective acoustic Planck constants, $\hbar_{\rm ac}$ and $/\!\!\!h_{\rm ac}$:
@@
The ground state of superfluid 4He represents the strongly interacting and strongly correlated ”quantum vacuum”. Two Planck mass scales $m_{4}s^{2}$ and $\hbar s/a$ are of the same order. That is why in this analogue of quantum vacuum the acoustic Planck constant $/\!\!\!h_{\rm ac}\sim a$, i.e. the length of the effective Planck constant is on the order of the effective Planck length (the interatomic distance).
@@ IX.9 Phase boundary between Universes with different $\hbar$
[…]
@@
[…]
@@ X Discussion
It is shown that the de Sitter Universe can be represented in terms of two dark components in equations (26)-(29). These are the dark energy (the energy of the vacuum) and the dark matter, which comes from the gravitational degrees of freedom. Their partial pressures compensate each other in the equilibrium de Sitter state. In this representation, the dark matter is responsible for the thermodynamics of the de Sitter state, which is characterized by the local temperature $T=H/\pi$ and is similar to the thermodynamics of the Zel’dovich stiff matter. Let us recall that the local temperature of the de Sitter state is twice the Gibbons-Hawking temperature related to the cosmological horizon, and the stiff matter behaviour of the dark matter component is the consequence of this value of the local temperature.
We tried to extend this two-fluid approach to the dynamics of the Universe, assuming that in dynamics the dark matter component is also equivalent to the Zel’dovich stiff matter. Then we obtained the power-law decay of both dark the matter component in Eq.(38) and the dark energy component in Eq.(37). While the state immediately after the Big Bang may have an energy density of the order of the Planck scale, $\epsilon\sim M^{4}_{\rm P}$, the energies of both components acquire the correct order of magnitude at the present time, $\epsilon\sim 10^{-120}M^{4}_{\rm P}$, thus offering a possible solution to the coincidence problem. Both components eventually decay to zero at $t\rightarrow\infty$ on the way to a full equilibrium attractor, the Minkowski quantum vacuum.
The reasonable decay law (37) for dark energy takes place only for the stiff matter. The question then arises: what might be the physical origin of this stiff matter and how is it related to the gravitational hard matter discussed in the Section IV. In principle, it is possible, that the stiff matter comes just from the gravitational dark matter, and the modified equations reflect the microscopic physics, which supports both the thermodynamics and dynamics of the de Sitter. The question then remains open as to how to describe the corresponding degrees of freedom that lead to the considered thermodynamics and dynamics.

### v10

@@ IX.8 Acoustic Planck constants
Returning to the two-fluid hydrodynamics, let us consider quasiparticles in Bose superfluid 4He. The flow of the liquid provides the acoustic gravity for ”relativistic” quasiparticles – phonons. The effective acoustic metric and thus the acoustic Planck constants, $\hbar_{\rm ac}$ and $h_{\rm ac}$, can be expressed in terms of the parameters of this liquid.
@@
$S_{\rm ph}=\frac{m_{4}}{2\hbar}\int d^{3}xdt\,n\left((\nabla\Phi)^{2}-\frac{1}{s^{2}}(\dot{\Phi}-{\bf v}\cdot\nabla\Phi)^{2}\right)=\,$
@@
The quantum mechanics of phonons demonstrates that it can be described by the effective acoustic Planck constants, $\hbar_{\rm ac}$ and $h_{\rm ac}$:
@@
The ground state of superfluid 4He represents the strongly interacting and strongly correlated ”quantum vacuum”. Two Planck mass scales $m_{4}s^{2}$ and $\hbar s/a$ are of the same order. That is why in this analogue of quantum vacuum the acoustic Planck constant $h_{\rm ac}\sim a$, i.e. the length of the effective Planck constant is on the order of the effective Planck length (the interatomic distance).
@@ IX.9 Phase boundary between Universes with different $\hbar$
The so-called multiple point principleNielsen1994 suggests that the quantum vacua are degenerate, and thus there is the boundary of the first order phase transition between our Universe and the Universe with very different properties.Nielsen2023 If so, it is not excluded that the two neighbouring Universes may have different Minkowski metric and thus different values of Planck parameters $\hbar$ and /h/\!\!\!h. If there is the thermal contact between these Universes, then in the thermal equilibrium one would have the following condition for temperatures and Planck constants:
@@
Another interesting question is what happens when $e_{\mu}^{a}\rightarrow 0$, i.e. we approach the critical point at which the interval $ds$ becomes identically zero and thus the spacetime disappears. According to Eq.(83, in this limit the both Planck constants approach infinity, $\hbar\rightarrow\infty$ and $h\rightarrow\infty$. This demonstrates the quantum correlations between the points separated by arbitrary large distances. However, there may be different limiting cases corresponding to different relations between $\hbar$, /h/\!\!\!h and $G$ in the Bronstein $cGh$ cube,Bronstein1933 ; Duff2002 see Ref. Ecker2025 and references therein.
@@ X Discussion
It is shown that the de Sitter Universe can be represented in terms of two dark components in equations (26)-(29). These are the dark energy (the energy of the vacuum) and the ”gravitational dark matter”, which comes from the gravitational degrees of freedom. The partial pressures of two components compensate each other in the equilibrium de Sitter state. In this representation, the gravitational dark matter is responsible for the thermodynamics of the de Sitter state, which is characterized by the local temperature $T=H/\pi$ and is similar to the thermodynamics of the Zel’dovich stiff matter. Let us recall that the local temperature of the de Sitter state is twice the Gibbons-Hawking temperature related to the cosmological horizon, and the stiff matter behaviour of the dark matter component is the consequence of this value of the local temperature.
We tried to extend this two-fluid approach to the dynamics of the Universe, assuming that in dynamics the dark matter component is also equivalent to the Zel’dovich stiff matter. Then we obtained the power-law decay of both dark components, the dark matter component in Eq.(38) and the dark energy component in Eq.(37). While the state immediately after the Big Bang may have an energy density of the order of the Planck scale, $\epsilon\sim M^{4}_{\rm P}$, the energies of both components acquire the correct order of magnitude at the present time, $\epsilon\sim 10^{-120}M^{4}_{\rm P}$, thus offering a possible route to solution of the coincidence problem. Both components eventually decay to zero at $t\rightarrow\infty$ on the way to a full equilibrium attractor, the Minkowski quantum vacuum.
The reasonable decay law (37) for dark energy is obtained in the model with the stiff matter. The question then arises: what might be the physical origin of this stiff matter and how is it related to the gravitational hard matter discussed in the Section IV. In principle, it is possible, that the stiff matter comes just from the gravitational dark matter, and the modified equations reflect the microscopic physics, which supports both the thermodynamics and dynamics of the de Sitter. The question then remains open as to how to describe the corresponding degrees of freedom that lead to the considered thermodynamics and dynamics.